solve elimination/addtion method
2c - 3d = 14
5c - 6d = 32
show steps

Answer by mananth(16946)

2c-3d=14 .............1
5c-6d=32 .............2
Eliminate y
multiply (1)by-2
Multiply (2) by 1
-4c+6d= -28
5c-6d=32
Add the two equations
c=4
plug value of c in (1)
2c-3d=14
8-3d=14
-3d=14 -8
-3d=6
d=-2